# Preference-Driven Multi-Objective Combinatorial Optimization with Conditional Computation

### v2

@@ 1 Introduction
To address these issues, we propose POCCO (Preference-driven multi-objective combinatorial Optimization with Conditional COmputation), a plug-and-play framework that augments neural MOCOP methods with two complementary mechanisms. First, POCCO introduces a conditional computation block into the decoder, where a sparse gating network dynamically routes each subproblem through either a selected subset of feed-forward (FF) experts or a parameter-free identity (ID) expert. This design enables subproblems to adaptively select computation routes (i.e., model structures) based on their context, efficiently scaling model capacity and facilitating more effective representation learning. Second, POCCO replaces raw scalarized rewards with pairwise preference learning. For each subproblem, the policy samples two trajectories, identifies the better one as the winner, and maximizes a Bradley–Terry (BT) likelihood based on the difference in their average log-likelihoods. Such comparative feedback guides the search toward policies that generate increasingly preferred solutions, enabling exploration of the most promising regions of the search space and more efficient convergence to higher-quality solutions. Our contributions are summarized as follows:
@@ 2.1 MOCOP
Definition 2 (Pareto Optimality). A solution $\pi^{*}\in\mathcal{X}$ is Pareto optimal if it is not dominated by any other solution. Accordingly, the Pareto set $\mathcal{P}$ is defined as all Pareto optimal solutions, i.e., $\mathcal{P}=\left\{\pi^{*}\in\mathcal{X}|\nexists\;\pi\in\mathcal{X}:\pi\prec\pi^{*}\right\}$. The Pareto front $\mathcal{F}$ is defined as images of Pareto optimal solutions in the objective space, i.e., =$\mathcal{F}=\left\{F(\pi^{*})|\pi^{*}\in\mathcal{P}\right\}$.
@@ 3.1 Overview
POCCO is a learning-based framework that trains a portfolio of policies to solve a set of scalarized subproblems $\{(\mathcal{G},\lambda_{i})\}_{i=1}^{N}$, obtained by decomposing an MOCOP instance. Instead of forcing a single policy to handle all subproblems which often yields bland and suboptimal behavior, POCCO promotes specialization: each policy is encouraged to focus on a subset of subproblems, yielding a diverse policy ensemble. Such diversity is known to enhance multi-task optimization [54] by expanding the exploration space and ultimately improving solution quality. Technically, we achieve this diversity by activating different subsets of model parameters through a CCO block, enabling distinct computational paths to emerge for different subproblems. Moreover, POCCO should encourage each policy to thoroughly explore the combinatorial solution space during training for reducing suboptimality. To achieve this, we replace raw rewards with preference signals. For each subproblem $(\mathcal{G},\lambda_{i})$, we construct a set of winning–losing solution pairs $\{(\pi^{w,j},\pi^{l,j})\}_{j=1}^{K}$. Training then proceeds by maximizing the likelihood of the winning solutions while minimizing that of the losing ones, following a BT-style objective. This preference-driven training encourages the learned policy to explore the most promising regions of the search space, leading to more efficient convergence toward higher-quality solutions. Notably, POCCO is a generic, plug-and-play framework that can seamlessly involve different neural solvers for MOCOPs. We demonstrate this by augmenting two SOTA methods, CNH [14] and WE-CA [6], in Section 4.
@@ 3.2 Conditional Computational Block
Formally, a CCO block consists of: 1) $m$ FF experts $\{E_{1},E_{2},\dots,E_{m}\}$ with independent trainable parameters; 2) a parameter-free identity expert $E_{m+1}$; 3) a router, implemented as a gating network $G$ parameterized by $W_{G}$, which determines how the inputs $\{h_{c}^{b}\}_{b=1}^{B}$ are routed to the experts; and 4) a skip connection followed by an instance normalization (IN) layer. Given a single context vector $h_{c}^{b}$, let $G(h_{c}^{b})\in\mathbb{R}^{m+1}$ denote the output of the gating network, which represents the expert selection probabilities, and let $E_{j}(h_{c}^{b})$ denote the output of the $j$-th expert. The output of the CCO block is:
@@
The sparse vector $G(h_{c}^{b})$ activates only a small subset of experts, either parameterized FF experts or the parameter-free ID expert, thereby enabling diverse computation paths while reducing computational overhead. A typical implementation uses a $\operatorname{Top}k$ operator that retains the $k$ largest logits and masks the rest with $\infty$- ∞. In this case, the gating network output is: $G(h_{c}^{b})=\operatorname{Softmax}\bigl{(}\operatorname{Top}k(h_{c}^{b}\cdot W_{G})\bigr{)}$.
Our proposed CCO block aligns with the principles of recent advances in efficiently scaling Transformer-based models along both width [49] and depth [46]. In specific, it combines of a mixture-of-experts (MoE) layer, implemented using multiple FF experts to widen the network, with a mixture-of-depths (MoD) layer, realized through an ID expert that allows inputs to skip computation. Within the CCO block, each subproblem is adaptively routed to only a small subset of experts, granting the model the expressiveness of a significantly wider network while preserving computational efficiency. As demonstrated in Section 4, this joint design achieves a better capacity–efficiency trade-off than scaling either dimension in isolation.
@@ 3.3 Preference-driven MOCO
Defining an implicit reward. Distinct from DRL training paradigms that rely on raw objective values, POCCO treats the average log-likelihood of a solution as an implicit reward $f_{\theta}$, directly relating preferences between solutions to their policy probabilities. This reward is inherently normalized by sequence length $|\pi|$, thereby mitigating length bias between winning and losing solution pairs.
@@ 4.1 Experimental settings
Training. We conduct extensive experiments to evaluate the effectiveness of the proposed POCCO across various MOCOPs, including multi-objective traveling salesman problem (MOTSP) [40] multi-objective capacitated vehicle routing problem (MOCVRP) [65], and multi-objective knapsack problem (MOKP) [23]. In MOTSP, the goal is to find a tour that visits all nodes exactly once, while minimizing multiple total path lengths, each computed based on a distinct set of coordinates associated with a specific objective. Regarding MOCVRP, a fleet of vehicles with limited capacity must serve all customer nodes and return to the depot, with the objectives of minimizing the total travel distance and the length of the longest individual route. As for MOKP, the problem involves selecting a subset of items, each with a weight and multiple objective-specific values. The objective is to maximize all objective values simultaneously, while ensuring the total weight stays within the knapsack capacity. In this work, we consider three commonly used problem sizes: $n=20/50/100$ for MOTSP and MOCVRP, and $n=50/100/200$ for MOKP.
@@
Baselines. We compare POCCO with a broad range of baseline methods across three categories, all employing weighted-sum (WS) scalarization to ensure fair comparison: (1) Single-model neural MOCO approaches: This includes PMOCO [39], and recent SOTA methods CNH [14], and WE-CA [6]. Both CNH and WE-CA are unified model trained across problem size $n\in\{20,21,\cdots,100\}$ (except $n\in\{50,51,\cdots,200\}$ for Bi-KP) (2) Multi-model neural MOCO approaches: This category covers methods like DRL-MOA [34], MDRL [68], and EMNH [7]. Specifically, DRL-MOA trains a separate POMO model for each of the $N$ subproblems, with the first model trained for 200 epochs and the rest fine-tuned for 5 epochs each using parameter transfer. MDRL and EMNH both initialize from a shared pretrained meta-model and fine-tune $N$ subproblem-specific models using the same network structure and training settings as in [7]. (3) Non-learnable approaches, including classical MOEAs and other problem-specific heuristics: MOEA/D [66] and NSGA-II [12], each run for 4,000 iterations, serve as representative decomposition-based and dominance-based MOEAs, respectively. MOCOP-specific MOEAs such as MOGLS [24], configured with 4,000 iterations and 100 local search steps per iteration, and PPLS/D-C [50], run for 200 iterations, are also considered. These methods use 2-opt heuristics for MOTSP and MOCVRP, and a greedy transformation heuristic [23] for MOKP. Finally, WS-LKH and WS-DP combine weighted-sum scalarization with powerful solvers, with LKH [20, 52] used for MOTSP and dynamic programming applied to MOKP.
Inference. We evaluate all methods using three metrics: average hypervolume (HV) [55], average gap, and total runtime per instance set. HV is a widely used indicator in multi-objective optimization that reflects both the convergence and diversity of the solution set. A higher HV indicates better performance. To ensure consistency, HV values are normalized to the range $[0,1]$ using the same reference point for all methods. The gap is defined as the relative difference between a method’s HV and the HV of POCCO-W. Methods with the “-Aug” suffix apply instance augmentation [39] to further improve performance. To evaluate statistical significance, we use the Wilcoxon rank-sum test [56] at a 1% significance level. The best result and others that are not significantly worse are marked in bold, while the second-best and statistically similar results are underlined. All experiments are implemented in Python and conducted on a machine with NVIDIA Ampere A100-80GB GPUs and an AMD EPYC 7742 CPU. The code and dataset will be released publicly upon acceptance.
@@ 4.2 Experimental results
Out-of-distribution size generalization analysis. We evaluate the generalization ability of the models on out-of-distribution, benchmark instances KroAB100/150/200 [41]. All neural methods are trained on Bi-TSP100, except for CNH, WE-CA, and POCCO, which are trained across varying sizes $n\in\{20,21,\dots,100\}$. The results, visualized in Fig. 3, show that POCCO-W-Aug consistently achieves the best generalization performance compared with other neural baselines. POCCO-C-Aug also outperforms CNH-Aug across all evaluated scenarios. Full benchmark results and experiments on larger-scale instances Bi-TSP150/200 are provided in Appendix H and Appendix I, respectively.
@@
Effectiveness of the PL. We assess the training efficiency of the PL by comparing it to REINFORCE on the WE-CA and POCCO-W models using the Bi-TSP100 dataset. As shown by the validation curves in Fig. 4(a), PL achieves faster convergence despite identical network architectures. Notably, for WE-CA, training with PL for 100 epochs reaches performance comparable to 200 epochs of REINFORCE. Similar improvements are observed for POCCO-W. These results demonstrate that PL effectively accelerates training process and achieves better performance with fewer training epochs.
Effectiveness of the CCO block. To evaluate the impact of the CCO block’s structure and placement, we compare POCCO-W with WE-CA and several POCCO variants: POCCO-E (CCO inserted in the encoder replacing the FF layer), POCCO-D (CCO replacing the final linear layer of MHA in the decoder, using MLP experts), POCCO-MOE (replacing CCO with a standard MoE layer), POCCO-MD (replacing CCO with three MoD layers), POCCO-MED (using MoE in the encoder and MoD in place of CCO), and POCCO-Emp (replacing the identity expert in CCO with an empty expert). As shown in Fig. 4(b), all variants outperform WE-CA on the in-distribution Bi-TSP100, with POCCO-W, POCCO-MOE, and POCCO-Emp achieving the most notable gains. On the out-of-distribution Bi-TSP200 in Fig. 4(c), only POCCO-W and POCCO-Emp maintain strong performance, while POCCO-MED performs worst, even underperforming WE-CA. These results highlight the importance of both the structure and placement of the CCO block for achieving strong generalization across in- and out-of-distribution settings.
Hyperparameter study. We conduct experiments to examine the impact of different key hyperparameters on POCCO’s performance. As detailed in Appendix J, the number of CCO block layers, the $\operatorname{Top}k$ value, and the temperature parameter $\beta$ all influence model effectiveness. For the problems studied, the desirable settings, as identified based on empirical results, are: one CCO block layer, $\operatorname{Top}k=2$, $\beta=3.5$ for bi-objective tasks, and $\beta=4.5$ for tri-objective tasks.

### v4

@@ 1 Introduction
To address these issues, we propose POCCO (Preference‑driven multi‑objective combinatorial Optimization with Conditional COmputation), a plug‑and‑play framework that augments neural MOCOP methods with two complementary mechanisms. First, POCCO introduces a conditional computation block into the decoder, where a sparse gating network dynamically routes each subproblem through either a selected subset of feed-forward (FF) experts or a parameter-free identity (ID) expert. This design enables subproblems to adaptively select computation routes (i.e., model structures) based on their context, efficiently scaling model capacity and facilitating more effective representation learning. Second, POCCO replaces raw scalarized rewards with pairwise preference learning. For each subproblem, the policy samples two trajectories, identifies the better one as the winner, and maximizes a Bradley–Terry (BT) likelihood based on the difference in their average log-likelihoods. Such comparative feedback guides the search toward policies that generate increasingly preferred solutions, enabling exploration of the most promising regions of the search space and more efficient convergence to higher-quality solutions. Our contributions are summarized as follows:
@@
2 Related Works
Traditional methods for MOCOP. MOCOPs are significantly harder to solve than their single‑objective counterparts. Classic cases such as the Traveling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP) are already NP‑hard, and adding multiple objectives only magnifies the difficulty [16]. Exact algorithms quickly become impractical for large instances or for problems that possess a vast Pareto set [18, 59]. Consequently, research has shifted toward heuristic methods that deliver high‑quality Pareto fronts within reasonable time limits [64, 25]. Among these, multi‑objective evolutionary algorithms (MOEAs) are widely used. They fall into two main categories: dominance‑based MOEAs [12, 11, 46] and decomposition‑based MOEAs [65, 15, 26]. Despite their popularity, MOEAs typically demand extensive manual design: practitioners must select and tune crossover, mutation, and selection operators along with many hyperparameters [63, 61, 50], and this labor‑intensive hand‑engineering often limits overall performance.
Preference optimization. Preference optimization methods, such as Direct Preference Optimization (DPO) [44] and Identity Preference Optimization (IPO) [2], have gained significant traction, particularly in large language model (LLM) training, due to their ability to directly optimize human preferences without relying on explicit reward modeling as required in reinforcement learning from human feedback (RLHF). These methods are typically designed for pairwise preference data, where human annotators identify a preferred output over a less preferred one in response to a given prompt. Another line of research explores simpler preference optimization objectives that do not depend on a reference model [21, 62]. Among them, SimPO [41] proposes using the average log-probability of a generated sequence as an implicit reward. This approach aligns more closely with the model’s generation process and improves computational and memory efficiency by eliminating the need for a separate reference model. Some studies [37] have explored applying preference optimization to SOCOPs, but its application to MOCOPs has been rarely investigated.
@@ 3.1 MOCOP
Definition 2 (Pareto Optimality). A solution $\pi^{*}\in\mathcal{X}$ is Pareto optimal if it is not dominated by any other solution. Accordingly, the Pareto set $\mathcal{P}$ is defined as all Pareto optimal solutions, i.e., $\mathcal{P}=\left\{\pi^{*}\in\mathcal{X}|\nexists\;\pi\in\mathcal{X}:\pi\prec\pi^{*}\right\}$. The Pareto front $\mathcal{F}$ is defined as images of Pareto optimal solutions in the objective space, i.e., $\mathcal{F}=\left\{F(\pi^{*})|\pi^{*}\in\mathcal{P}\right\}$.
@@ 4.1 Overview
POCCO is a learning-based framework that trains a portfolio of policies to solve a set of scalarized subproblems $\{(\mathcal{G},\lambda_{i})\}_{i=1}^{N}$, obtained by decomposing an MOCOP instance. Instead of forcing a single policy to handle all subproblems, which often yields bland and suboptimal behavior, POCCO promotes specialization: each policy is encouraged to focus on a subset of subproblems, yielding a diverse policy ensemble. Such diversity is known to enhance multi-task optimization [53] by expanding the exploration space and ultimately improving solution quality. Technically, we achieve this diversity by activating different subsets of model parameters through a CCO block, enabling distinct computational paths to emerge for different subproblems. Moreover, POCCO should encourage each policy to thoroughly explore the combinatorial solution space during training for reducing suboptimality. To achieve this, we replace raw rewards with preference signals. For each subproblem $(\mathcal{G},\lambda_{i})$, we construct a set of winning–losing solution pairs $\{(\pi^{w,j},\pi^{l,j})\}_{j=1}^{K}$. Training then proceeds by maximizing the likelihood of the winning solutions while minimizing that of the losing ones, following a BT-style objective. This preference-driven training encourages the learned policy to explore the most promising regions of the search space, leading to more efficient convergence toward higher-quality solutions. Notably, POCCO is a generic, plug-and-play framework that can seamlessly involve different neural solvers for MOCOPs. We demonstrate this by augmenting two SOTA methods, CNH [14] and WE-CA [6], in Section 5.
@@ 4.2 Conditional Computational Block
Formally, a CCO block consists of: 1) $m$ FF experts $\{E_{1},E_{2},\dots,E_{m}\}$ with independent trainable parameters; 2) a parameter-free identity expert $E_{m+1}$; 3) a router, implemented as a gating network $G$ parameterized by $W_{G}$, which determines how the inputs $\{h_{c}^{b}\}_{b=1}^{B}$ are routed to the experts; and 4) a skip connection followed by an instance normalization (IN) layer. Given a single context vector $h_{c}^{b}$, let $G(h_{c}^{b})\in\mathbb{R}^{m+1}$ denote the output of the gating network, which represents the expert selection probabilities, and let $E_{j}(h_{c}^{b})$ denote the output of the jj‑th expert. The output of the CCO block is:
@@
The sparse vector $G(h_{c}^{b})$ activates only a small subset of experts, either parameterized FF experts or the parameter‑free ID expert, thereby enabling diverse computation paths while reducing computational overhead. A typical implementation uses a $\operatorname{Top}k$ operator that retains the $k$ largest logits and masks the rest with -$\infty$. In this case, the gating network output is: $G(h_{c}^{b})=\operatorname{Softmax}\bigl(\operatorname{Top}k(h_{c}^{b}\cdot W_{G})\bigr)$.
Our proposed CCO block aligns with the principles of recent advances in efficiently scaling Transformer-based models along both width [48] and depth [45]. In specific, it combines a mixture‑of‑experts (MoE) layer, implemented using multiple FF experts to widen the network, with a mixture‑of‑depths (MoD) layer, realized through an ID expert that allows inputs to skip computation. Within the CCO block, each subproblem is adaptively routed to only a small subset of experts, granting the model the expressiveness of a significantly wider network while preserving computational efficiency. As demonstrated in Section 5, this joint design achieves a better capacity–efficiency trade-off than scaling either dimension in isolation.
@@ 4.3 Preference-driven MOCO
Defining an implicit reward. Distinct from DRL training paradigms that rely on raw objective values, POCCO treats the average log‑likelihood of a solution as an implicit reward $f_{\theta}$, directly relating preferences between solutions to their policy probabilities. This reward is inherently normalized by sequence length $|\pi|$, thereby mitigating length bias between winning and losing solution pairs.
@@ 5.1 Experimental settings
Training. We conduct extensive experiments to evaluate the effectiveness of the proposed POCCO across various MOCOPs, including multi-objective traveling salesman problem (MOTSP) [39], multi-objective capacitated vehicle routing problem (MOCVRP) [64], and multi-objective knapsack problem (MOKP) [23]. In MOTSP, the goal is to find a tour that visits all nodes exactly once, while minimizing multiple total path lengths, each computed based on a distinct set of coordinates associated with a specific objective. Regarding MOCVRP, a fleet of vehicles with limited capacity must serve all customer nodes and return to the depot, with the objectives of minimizing the total travel distance and the length of the longest individual route. As for MOKP, the problem involves selecting a subset of items, each with a weight and multiple objective-specific values. The objective is to maximize all objective values simultaneously, while ensuring the total weight stays within the knapsack capacity. In this work, we consider three commonly used problem sizes: $n=20/50/100$ for MOTSP and MOCVRP, and $n=50/100/200$ for MOKP.
@@
Baselines. We compare POCCO with a broad range of baseline methods across three categories, all employing weighted-sum (WS) scalarization to ensure fair comparison: (1) Single-model neural MOCO approaches: This includes PMOCO [38], and recent SOTA methods CNH [14], and WE-CA [6]. Both CNH and WE-CA are unified models trained across problem size $n\in\{20,21,\cdots,100\}$ (except $n\in\{50,51,\cdots,200\}$ for Bi-KP) (2) Multi-model neural MOCO approaches: This category covers methods like DRL-MOA [34], MDRL [67], and EMNH [7]. Specifically, DRL-MOA trains a separate POMO model for each of the $N$ subproblems, with the first model trained for 200 epochs and the rest fine-tuned for 5 epochs each using parameter transfer. MDRL and EMNH both initialize from a shared pretrained meta-model and fine-tune $N$ subproblem-specific models using the same network structure and training settings as in [7]. (3) Non-learnable approaches, including classical MOEAs and other problem-specific heuristics: MOEA/D [65] and NSGA-II [12], each run for 4,000 iterations, serve as representative decomposition-based and dominance-based MOEAs, respectively. MOCOP-specific MOEAs such as MOGLS [24], configured with 4,000 iterations and 100 local search steps per iteration, and PPLS/D-C [49], run for 200 iterations, are also considered. These methods use 2-opt heuristics for MOTSP and MOCVRP, and a greedy transformation heuristic [23] for MOKP. Finally, WS-LKH and WS-DP combine weighted-sum scalarization with powerful solvers, with LKH [20, 51] used for MOTSP and dynamic programming applied to MOKP.
Inference. We evaluate all methods using three metrics: average hypervolume (HV) [54], average gap, and total runtime per instance set. HV is a widely used indicator in multi-objective optimization that reflects both the convergence and diversity of the solution set. A higher HV indicates better performance. To ensure consistency, HV values are normalized to the range $[0,1]$ using the same reference point for all methods. The gap is defined as the relative difference between a method’s HV and the HV of POCCO-W. Methods with the “-Aug” suffix apply instance augmentation [38] to further improve performance. To evaluate statistical significance, we use the Wilcoxon rank-sum test [55] at a 1% significance level. The best results and others that are not significantly worse are marked in bold, while the second-best and statistically similar results are underlined. All experiments are implemented in Python and conducted on a machine with NVIDIA Ampere A100-80GB GPUs and an AMD EPYC 7742 CPU. The code and dataset are publicly released for reproducibility. [^1] https://github.com/mingfan321/POCCO
@@ 5.2 Experimental results
Out-of-distribution size generalization analysis. We evaluate the generalization ability of the models on out-of-distribution, benchmark instances KroAB100/150/200 [40]. All neural methods are trained on Bi-TSP100, except for CNH, WE-CA, and POCCO, which are trained across varying sizes $n\in\{20,21,\dots,100\}$. The results, visualized in Fig. 3, show that POCCO-W-Aug consistently achieves the best generalization performance compared with other neural baselines. POCCO-C-Aug also outperforms CNH-Aug across all evaluated scenarios. Full benchmark results and experiments on larger-scale instances Bi-TSP150/200 are provided in Appendix I and Appendix J, respectively.
@@
Effectiveness of PL. We assess the training efficiency of PL by comparing it to REINFORCE on the WE-CA and POCCO-W models using the Bi-TSP100 dataset. As shown by the validation curves in Fig. 4(a), PL achieves faster convergence despite identical network architectures. Notably, for WE-CA, training with PL for 100 epochs reaches performance comparable to 200 epochs of REINFORCE. Similar improvements are observed for POCCO-W. These results demonstrate that PL accelerates the training process and achieves better performance with fewer training epochs.
Effectiveness of the CCO block. To evaluate the impact of the CCO block’s structure and placement, we compare POCCO-W with WE-CA and several POCCO variants: POCCO-E (CCO inserted in the encoder replacing the FF layer), POCCO-D (CCO replacing the final linear layer of MHA in the decoder, using MLP experts), POCCO-ME (replacing CCO with a standard MoE layer), POCCO-MD (replacing CCO with three MoD layers), POCCO-MED (using MoE in the encoder and MoD in place of CCO in the decoder), and POCCO-Emp (replacing the identity expert in CCO with an empty expert). As shown in Fig. 4(b), all variants outperform WE-CA on the in-distribution Bi-TSP100, with POCCO-W, POCCO-ME, and POCCO-Emp achieving the most notable gains. On the out-of-distribution Bi-TSP200 in Fig. 4(c), only POCCO-W and POCCO-Emp maintain strong performance, while POCCO-ME performs worst, even underperforming WE-CA. These results highlight the importance of both the structure and placement of the CCO block for achieving strong generalization across in- and out-of-distribution settings.
Hyperparameter study. We conduct experiments to examine the impact of different key hyperparameters on POCCO’s performance. As detailed in Appendix L, the number of CCO block layers, the number of experts, the $\operatorname{Top}k$ value, and the temperature parameter $\beta$ all influence model effectiveness. For the problems studied, the desirable settings, as identified based on empirical results, are: one CCO block layer, four FF experts and one ID expert, $\operatorname{Top}k=2$, $\beta=3.5$ for bi-objective tasks, and $\beta=4.5$ for tri-objective tasks.